# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v1

@@ 4. Weakly-supervised Video Anomaly Detection
Inspired by multiple instance learning (MIL) (Herrera et al., 2016), WsVAD organizes videos into bags containing several instances. All segment instances from the same bag share a video-level label. Here, indicates negative bags while indicates positive ones containing at least one anomalous instance. WsVAD strikes a balance between performance and data preparation cost, bridging the gap between UVAD with unreliable results and supervised learning with fine labels. Specifically, unlike unsupervised methods that train models only on regular events and simplely consider all unseen samples as anomalies, WsVAD is trained with positive samples, enabling it to effectively understand the inherent differences between normal and anomalous instances. Consequently, existing research indicates that WsVAD yields more reliable results and outperforms UVAD schemes with lower false alarm rates in real-world tasks, such as crime behavior identification and traffic accident detection. Annotating long video sequences frame by frame for supervised learning is often impractical. For instance, the training set of UCF-Crime dataset includes 1,610 long videos containing 13,741,393 frames in total. In contrast, WsVAD only requires coarse semantic video-level annotations, with each video needing only a discrete binary label. Various types of anomalies are roughly labeled as 1, requiring minimal expert involvement in data annotation and significantly reducing labor costs compared to fine-grained labeling. Thus, although WsVAD no longer adheres to the open-world task setting of the AD community and can only detect specific anomalies, it has garnered widespread attention in recent years due to its stable performance and reliable results. Subsequent works typically adopt the MIL regression task setting proposed by Sultani et al. (Sultani et al., 2018) and utilize their concurrently open-sourced UCF-Crime dataset as a benchmark. To further validate the generalization ability of weakly-supervised models, some researchers relocate positive samples from the test set of UVAD datasets to the training sets and provide video-level labels, proposing reconfigured datasets such as UCSD Ped2 and ShanghaiTech Weakly for WsVAD validation.
@@ 4.1. Taxonomy and Advances
The superior performance of WsVAD on real-world videos has inspired researchers to develop VAD models tailored to the complexities of real-life scenarios. In 2020, Wu et al. (Wu et al., 2020) extended VAD research from single-modal video pattern analysis to multi-modal learning to leverage heterogeneous data in real-world scenes for enhanced anomaly event detection. Their collected VD-violence dataset is the first multi-modal VAD benchmark, comprising RGB images and audio modalities, focusing on violence behavior detection in complex scenes. This work not only provides the first multi-modal VAD dataset and solution but also motivates researchers in the community to explore anomaly clues from multi-modal data such as audio and text, leading to a new wave of VAD development: multi-modal WsVAD. In this section, we comprehensively review the latest advancements in both single-modal and multi-modal WsVAD, and illustrate the foundational knowledge and specific implementations required for WsVAD using the Multiple Instance Regression (MIR) framework and the Local-Global Network (LGN) as examples.
@@ 4.1.1. Unimodal Methods
Unimodal WsVAD (UWsVAD) methods focus on extracting anomaly-related cues from RGB image sequences. These methods generally follow a three-step process: 1) preprocessing videos into several non-overlapping segments, 2) extracting spatial-temporal features using models like Convolutional 3D Networks (Tran et al., 2015) and Inflated 3D Networks (Carreira and Zisserman, 2017), and 3) computing anomaly scores using a multi-instance ranking loss to differentiate between normal and anomalous instances. The Multi-Instance Ranking (MIR) framework (Sultani et al., 2018) first introduced MIL to WsVAD with a focus on predicting higher anomaly scores for anomalous segments while minimizing score fluctuations. For more detailed description of the existing UWsVAD methods, please refer to Section 1.2.1 of Supplement.
4.1.2. Multi-modal Methods
Multi-modal approaches in WsVAD combine video with audio and other data types. These methods face challenges due to limited benchmark datasets and comparison standards, which has hindered widespread adoption. A key milestone in this domain is HL-Net (Wu et al., 2020), introduced in 2020, marking a significant contribution to multi-modal VAD research with the open-sourced XD-Violence dataset. This dataset has spurred further research by providing a benckmarks for comparisons. Existing multi-modal research mainly explores the fusion of video, audio, and optical flow, as stated in Section 1.2.2 of Supplement.
@@ 4.2. Classic WsVAD Models
We select two representative methods from unimodal and multimodal WsVAD, i.e., Multi-Instance Ranking (MIR) (Sultani et al., 2018) and Holistic and Localized Network (HL-Net) (Wu et al., 2020), to elaborate on the concepts of MIL and multimodal information processing in WsVAD research, respectively. Specifically, MIR has laid the foundation for the MIL-based solution, marking a milestone in VAD research. However, its optimization objectives and detailed implementation are challenging for many researchers in the AD and CV communities to understand. Therefore, we provide a comprehensive exposition to the motivation, theoretical logic, and related knowledge. In contrast, HL-Net is the first model for multimodal violence detection. The simultaneous released XD-violence dataset has inspired lots of researchers from the AD and multimodal understanding fields to delve into the emerging hotspot of multimodal VAD. We choose this method as a case study to disseminate knowledge of multimodal understanding to the AD community, aiming to propel multimodal VAD from simple-modal fusion towards systematic multimodal anomaly clues exploration, please see Section 2.2 of Supplement for more information.
@@ 6. Model Evaluation
In this section, we elaborate on the characteristics of prevailing datasets and common evaluation metrics. Based on the annotations, we categorize existing datasets into unsupervised, weakly supervised, and supervised, as presented in Table 3. Some datasets for emerging WsVAD and FuVAD methods are reconstructed from UVAD, so we provide explanations in corresponding sections, refraining from separate introductions. Additionally, we provide the first systematic WSVAD evaluation system, categorizing existing metrics into accuracy-oriented and cost-oriented.
6.1. Public Datasets
6.1.1. Unsupervised Datasets
Subway (Adam et al., 2008) focuses on abnormal event detection in subway stations, comprising two subsets: Entrance and Exit, with abnormal behaviors depicted as shown in Fig. 8. The entrance is used to detect fare evasion behaviors like jumping over turnstiles. In contrast, abnormal events in the Exit include 1) appearance anomalies, such as cleaners wearing attire different from regular passengers, and 2) motion anomalies, such as people walking in the opposite direction of the exit. This dataset was collected 20 years ago, so its resolution is low. Moreover, due to overly coarse annotations and lack of clear rules defining abnormal events, recent VAD works have discontinued using this dataset. Therefore, we recommend researchers to use larger and more clearly annotated datasets, such as UCSD Ped2, CUHK Avenue, and ShanghaiTech, for validation.
UMN (Cui et al., 2011) consists of 11 long video sequences captured from 3 different scenes. Unlike Subway, this dataset includes behaviors pre-set by the collectors and volunteers acting them out, as shown in Fig. 8. Abnormal events are defined as sudden evacuations caused by panic in crowds, simulating mass anomalies in the real world due to terror attacks and fires. Due to the lack of convincing application scenarios, neglect of diversity, and low frequency of real-world anomalies, this dataset is no longer used in UVAD.
UCSD Pedestrian (Li et al., 2013) is one of the most widely used UVAD benchmarks collected by researchers from the campus of University of California, San Diego. Due to its realistic scenes and clear definition of abnormal events, it is widely used to evaluate models’ potential applications in public safety. According to camera angles, UCSD Pedestrian is divided into Ped1 and Ped2 subsets. The latter is more favored by existing methods because the camera view is parallel to the road direction, eliminating the problem of changing object sizes with spatial changes. Recent methods have achieved remarkable frame-level AUCs exceeding 98% on UCSD Ped2, while the best performance on Ped1 is still below 94%, indicating that performance on Ped2 has saturated. Considering that cross-view videos are inevitable in the real world, we recommend that future researchers consider using Ped1 for validating and empowering UVAD models to overcome perspective issues. Some researchers move abnormal samples from the test set of UCSD Ped2 to the training set to meet the requirements of weakly supervised training, proposing a reconstructed UCSD Ped2 dataset for WsVAD testing.
CUHK Avenue (Lu et al., 2013) is similar to UCSD Pedestrian in scene selection and data structures, with only the test set containing 47 abnormal events captured from a university campus. However, CUHK Avenue considers more realistic situations. Collectors collected appearance-only, motion-only, and appearance-motion united anomalies to comprehensively simulate the diversity and complexity of real-world anomalies. Therefore, the appearance and motion patterns of videos in CUHK Avenue are more complex, making it one of the most challenging UVAD benchmarks. Additionally, CUHK Avenue inspired later researchers to collect UVAD datasets by simulating abnormal behaviors, such as ShanghaiTech (Liu et al., 2018b) and NWPU Campus (Cao et al., 2023).
ShanghaiTech (Liu et al., 2018b) breaks the single-scene limitations of UCSD Ped2 and CUHK Avenue, collecting 158 abnormal events from 13 scenes. Compared to the above datasets, it has a larger data scale and more complex anomaly types. Therefore, the performance of existing UVAD methods on ShanghaiTech is generally lower (AUC ), indicating the need for further exploration in cross-scene UVAD (Aich et al., 2023). Moreover, researchers proposed ShanghaiTech Weakly for WsVAD. Recent studies have shown that with similar scenarios and anomalies, the performance of WsVAD methods on the re-organized dataset significantly outperforms UVAD models on the original one. For example, mainstream WsVAD models like MIST (Feng et al., 2021a) and RTFM (Tian et al., 2021) achieve a frame-level AUC of up to 98%, surpassing state-of-the-art UVAD by nearly 20%. This phenomenon indicates that weakly supervised methods are more applicable in real-world scenarios.
Street Scene (Ramachandra and Jones, 2020) dataset is the first UVAD benchmark dedicated to anomaly detection in complex traffic, suitable for validating models’ deployment potential in smart city applications. The authors used static USB cameras to collect real street scenes (sidewalks, bike lanes, and roadways) in Cambridge, Massachusetts, USA, covering various traffic participants like pedestrians, non-motorized vehicles, and motor vehicles. Following the UVAD task setting, all 46 training videos depict normal events, while 35 test samples contain 205 abnormalities categorized into 17 classes. This dataset contains rich traffic behaviors, such as vehicle-centric driving, turning, and stopping, pedestrian-centric walking, jogging, and pushing strollers, and non-motorized vehicle-centric biking. Additionally, due to daylight angle changes, handling this dataset also requires addressing the effects of building shadows and moving backgrounds in urban scenes. Apart from temporal annotations, collectors also provided trajectory-level annotations, i.e., bounding boxes and IDs marking the boundaries of anomalies for fine-grained spatial localization learning.
NWPU Campus (Cao et al., 2023) is currently the largest UVAD dataset, containing 547 videos collected from 43 scenes, totaling 76.6GB. As shown in Table 3, this dataset’s data scale, scene quantity, and abnormal class diversity exceed those of existing unsupervised datasets, allowing for thorough testing of model effectiveness and generalization. It’s worth noting that the collectors considered the scene relevance of abnormal events, meaning similar behavioral patterns might be classified as different categories in different scenes. Therefore, during data collection, they simulated such behaviors extensively to enhance the dataset’s consistency with real-world scenes.
6.1.2. Weakly Supervised Datasets
UCF-Crime (Sultani et al., 2018) established the first benchmark for WsVAD and continuously promoted weakly supervised methods based on MIL. This dataset collected 1,900 unedited surveillance videos from the internet containing 850 crime events such as theft, robbery, and road accidents. Collectors categorized these anomalies into 13 classes, labeling videos containing such events as 1, while others were labeled as 0. This weak semantic annotation, termed video-level labels, is used to supervise model learning of finely detailed frame-level labels with solid semantics. Unlike UVAD, which only uses regular videos for model training, the UCF-Crime training set contains positive and negative samples, learning the differences between normal and abnormal by contrasting their pattern differences. We display abnormal samples from the UCF-Crime dataset in Fig. 8, including road accidents, shoplifting, stealing, shooting, and robbery. The anomalies defined in WsVAD align more with human recognition and have a higher degree of realism. Since predefined abnormal events are seen during the training phase, WsVAD exhibits more stable performance in detecting specific category anomalies, with less susceptibility to scene changes and external environmental influences. Subsequent weakly supervised datasets, such as XD-Violence (Wu et al., 2020) and these reconstructed from UVAD datasets like UCSD Ped2 and ShanghaiTech, also follow the data settings of UCF-Crime, with both training and testing sets containing positive and negative samples. However, the UCF-Crime dataset assumes VAD is a closed-set problem, as abnormal events in the test set have all appeared in the training set, thus unable to verify the model’s ability to discriminate anomalies in an open world.
XD-Violence (Wu et al., 2020) is the first mainstream large-scale VAD benchmark with multimodal data, including videos and synchronized audio. The authors collected various violent behaviors endangering personal safety from the internet as target anomalies, including riots, aggressions, shootings, and traffic accidents, as shown in Figure 8. Besides real surveillance videos from the internet, this dataset also includes audiovisual files from games and movies, significantly increasing the detection difficulty. XD-Violence opens a new avenue for VAD research, i.e., mining anomaly-related clues from heterogeneous data. Consistent with UCF-Crime (Sultani et al., 2018), XD-Violence provides coarse semantic video-level annotations. Each video only requires one label without annotating specific categories of anomalies, making the data preparation cost acceptable. Since the collectors did not consider the open-set characteristics of AD tasks, the anomaly event categories in the training and testing sets are consistent. Therefore, this dataset can only be used to test closed-set models. We recommend that future researchers place unseen anomalies in the testing set to develop multimodal VAD models with open-set detection capabilities.
6.1.3. Supervised Datasets
The high cost of collecting anomalies and fine-grained annotations makes existing VAD research typically avoid supervised learning requiring frame-level or pixel-level supervision. However, the rise of virtual data engines has made it feasible to construct diverse abnormal events and synchronously obtain precise labels. To this end, Acsintoae et al. (Acsintoae et al., 2022) constructed the first supervised VAD dataset named UBnormal by simulating virtual anomalies, providing a reference for addressing the data and label preparation challenges with virtual engines. In addition, this dataset can address the close-set concern of WsVAD tasks. The authors set up a validation set, including positive samples of different categories from the training set, to encourage models to detect unseen anomalies.
6.2. Evaluation Metrics
@@
The existing works typically evaluate the proposed method from two perspectives: detection accuracy and model cost. On the one hand, models are expected to detect noteworthy anomalies as accurately as possible. Considering the varying influences of false positives and false negatives, along with the highly imbalanced data, researchers have proposed multiple quantitative metrics such as Area Under the Receiver Operating Characteristic curve (AUROC), Area Under Precision-Recall curve (AUPR), and Detection Rate, to assess model accuracy. Additionally, anomaly score curves are commonly used to qualitatively demonstrate the sensitivity to abnormal intervals, while prediction error maps are widely employed in reconstruction-based UVAD methods to visualize the performance of spatial localization. On the other hand, metrics like parameter size and inference speed determine whether the model can be deployed on resource-constrained devices. Thus, we categorize existing metrics based on the orientation and evaluation dimension, as illustrated in Fig. 9.
@@
Due to the highly imbalanced nature of positive and negative samples in NSVAD, some common evaluation metrics for classification tasks, such as accuracy, are not applicable. For example, a model biased towards outputting label 0 thus missing anomalous events, would be incorrectly evaluated as good under such metrics. Using a single threshold to simply assess a model’s ability to differentiate between normal and abnormal patterns is not wise. For instance, with a threshold of 0.5, a model that consistently outputs scores slightly below 0.5 for regular events and slightly above 0.5 for anomalies would be considered perfect because it would show optimal performance across various metrics. However, such a model may not have learned the inherent differences between normal and abnormal patterns well, resulting in minimal score gaps between the two, which could lead to failure in detecting subtle anomalies and normal instances with data bias in complex scenarios. Therefore, researchers have introduced Receiver Operating Characteristic (ROC) curves, which measure VAD models more comprehensively by selecting multiple thresholds. Specifically, this curve plots the True Negative Rate (TNR) against the True Positive Rate (TPR) at various thresholds. The Area Under the ROC curve (AUROC) has been the most widely used VAD evaluation metric. An ideal AUROC value of 1 indicates a model that outputs a score of 0 for all negative samples and 1 for all positive samples, aligning with our expectations. Considering that TN usually exceeds TP, researchers argue that Average Precision, i.e., the Area Under the Precision-Recall (AUPR) curve, is more suitable for evaluating anomaly detection tasks. The PR curve depicts precision and recall (i.e., TPR) at specific thresholds. The point on this curve where Precision equals Recall is the balance point. Currently, multi-modal anomaly detection models primarily use AP for quantitative evaluation. Anomaly score curve is commonly used to intuitively demonstrate the model’s response to anomalous events, presenting the temporal localization capability. In contrast, prediction error maps are often used to assess UVAD model’s spatial localization capability.

### v2

@@ 4. Weakly-supervised Video Anomaly Detection
Inspired by multiple instance learning (MIL) (Herrera et al., 2016), WsVAD organizes videos into bags containing several instances. All segment instances from the same bag share a video-level label (Huang et al., 2024a). Here, indicates negative bags while indicates positive ones containing at least one anomalous instance. WsVAD strikes a balance between performance and data preparation cost, bridging the gap between UVAD with unreliable results and supervised learning with fine labels. Specifically, unlike unsupervised methods that train models only on regular events and simply consider all unseen samples as anomalies, WsVAD is trained with positive samples, enabling it to effectively understand the inherent differences between normal and anomalous instances. Consequently, existing research indicates that WsVAD yields more reliable results and outperforms UVAD schemes with lower false alarm rates in real-world tasks, such as crime behavior identification and traffic accident detection (Liu et al., 2023d). Annotating long video sequences frame by frame for supervised learning is often impractical. For instance, the training set of the UCF-Crime dataset includes 1,610 long videos containing 13,741,393 frames in total. In contrast, WsVAD only requires coarse semantic video-level annotations, with each video needing only a discrete binary label. Various types of anomalies are roughly labeled as 1, requiring minimal expert involvement in data annotation and significantly reducing labor costs compared to fine-grained labeling. Thus, although WsVAD no longer adheres to the open-world task setting of the AD community and can only detect specific anomalies, it has garnered widespread attention in recent years due to its stable performance and reliable results. Subsequent works typically adopt the MIL regression task setting proposed by Sultani et al. (Sultani et al., 2018) and utilize their concurrently open-sourced UCF-Crime dataset as a benchmark. To further validate the generalization ability of weakly-supervised models, some researchers relocate positive samples from the test set of UVAD datasets to the training sets and provide video-level labels, proposing reconfigured datasets such as UCSD Ped2 and ShanghaiTech Weakly for WsVAD validation.
@@ 4.1. Taxonomy and Advances
The superior performance of WsVAD on real-world videos has inspired researchers to develop VAD models tailored to the complexities of real-life scenarios. In 2020, Wu et al. (Wu et al., 2020) extended VAD research from single-modal video pattern analysis to multimodal learning to leverage heterogeneous data in real-world scenes for enhanced anomaly event detection. Their collected XD-violence dataset is the first multimodal VAD benchmark, comprising RGB images and audio modalities, focusing on violence behavior detection in complex scenes. This work not only provides the first multimodal VAD dataset and solution but also motivates researchers in the community to explore anomaly clues from multimodal data such as audio and text, leading to a new wave of VAD development: multimodal WsVAD. In this section, we comprehensively review the latest advancements in both single-modal and multimodal WsVAD, and illustrate the foundational knowledge and specific implementations required for WsVAD using the Multiple Instance Regression (MIR) framework and the Local-Global Network (HL-Net) as examples.
Unimodal methods take only visual data as input and attempt to learn the pattern differences between normal and anomalous events based on the appearance and motion semantics reflected in the RGB sequences. In contrast, multimodal methods explore additional data modalities beyond visuals, such as audio and text, using them as complementary semantics to improve the model’s anomaly detection capabilities. Although the inclusion of additional modalities increases the data processing and model training costs, existing studies show that audio and text can significantly enhance WsVAD’s ability to detect anomalies, especially in cases where visual information alone cannot distinguish anomalies effectively. Since multimodal methods still need to model spatiotemporal patterns in the visual modality and continue to use the same weak supervision task setup and multi-instance learning strategies as unimodal approaches, certain strategies already explored and validated in unimodal WsVAD—such as dataset bias correction, label noise reduction, and hinge loss optimization—can be incorporated into multimodal WsVAD methods, particularly for the visual data processing branch.
@@ 4.1.1. Unimodal Methods
Unimodal WsVAD (UWsVAD) methods focus on extracting anomaly-related cues from RGB image sequences (Liu et al., 2022d). These methods generally follow a three-step process: 1) preprocessing videos into several non-overlapping segments, 2) extracting spatial-temporal features using models like Convolutional 3D Networks (Tran et al., 2015) and Inflated 3D Networks (Carreira and Zisserman, 2017), and 3) computing anomaly scores using a multi-instance ranking loss to differentiate between normal and anomalous instances. The Multi-Instance Ranking (MIR) framework (Sultani et al., 2018) first introduced MIL to WsVAD with a focus on predicting higher anomaly scores for anomalous segments while minimizing score fluctuations. For a more detailed description of the existing UWsVAD methods, please refer to Section 1.2.1 of Supplement. Moreover, given that VsVAD datasets are typically collected from real-world scenarios and contain identity-sensitive information such as faces, Fioresi et al. (Fioresi et al., 2023) proposed a privacy-preserving VAD framework named TeD-SPAD. They first anonymized video frames using a UNet to eliminate privacy information before using the I3D network to extract spatial-temporal features. Results showed that TeD-SPAD successfully prevented 32% of visual information leakage.
4.1.2. Multimodal Methods
Multimodal approaches in WsVAD integrate various data types, primarily focusing on the fusion of video with audio (Wei et al., 2022a) and text (Wu et al., 2024b). These methods face significant challenges due to limited benchmark datasets and standard comparison metrics, which have hindered their widespread adoption.
One of the hallmark advancements in multimodal VAD is the effective fusion of video and audio data. Existing research (Yu et al., 2022; Pu and Wu, 2022; Wei et al., 2022b) primarily uses the XD-Violence (Wu et al., 2020) dataset as the evaluation benchmark, which introduces the audio into video violence detection. More information about the video-audio-based VAD methods is provided in Section 1.2.2 of Supplement.
Moreover, with the rise of Visual-Language Learning (VLL), particularly through the emergence of Large Language Models (LLMs), researchers have begun to leverage textual information to further enhance VAD models’ performance. Pre-trained Vision-Language Large Models (VLLMs) can describe appearance and motion information in videos without prior samples, embedding such text as prompts into visual representations. This integration represents a significant advancement in the fusion of video and text, which improves the model’s ability to express complex anomalies. The incorporation of textual information not only enhances generalizability and interpretability but also gives rise to new research tasks of practical value, such as video anomaly retrieval (Wu et al., 2024b) and open vocabulary VAD (Wu et al., 2024c), as illustrated in Sections 2.4 and 2.5.
In surveillance videos, it is often difficult to capture synchronized audio or text data, making unimodal methods the mainstream approach. However, with the rise of live streaming platforms and the film industry, multimodal methods will play a crucial role in online video moderation and content detection for TV shows, films, and animations. On the one hand, the context of such content is often highly complex and varied, with anomalies taking many forms, such as visual violence or non-compliant audio or text, making it difficult for unimodal methods relying solely on RGB sequences to handle effectively. On the other hand, these types of videos typically capture synchronized audio and provide subtitles or other multimodal data, allowing multimodal methods to function without additional data preparation costs.
@@ 4.2. Classic WsVAD Models
We select two representative methods from unimodal and multimodal WsVAD, i.e., Multi-Instance Ranking (MIR) (Sultani et al., 2018) and Holistic and Localized Network (HL-Net) (Wu et al., 2020), to elaborate on the concepts of MIL and multimodal information processing in WsVAD research, respectively. Specifically, MIR has laid the foundation for the MIL-based solution, marking a milestone in VAD research. However, its optimization objectives and detailed implementation are challenging for many researchers in the AD and CV communities to understand. Therefore, we provide a comprehensive exposition of the motivation, theoretical logic, and related knowledge. In contrast, HL-Net is the first model for multimodal violence detection. The simultaneously released XD-violence dataset has inspired lots of researchers from the AD and multimodal understanding fields to delve into the emerging hotspot of multimodal VAD. We choose this method as a case study to disseminate knowledge of multimodal understanding to the AD community, aiming to propel multimodal VAD from simple-modal fusion towards systematic multimodal anomaly clues exploration, please see Section 2.2 of Supplement for more information.
@@ 6. Model Evaluation
In this section, we elaborate on the characteristics of prevailing datasets and common evaluation metrics. Based on the annotations, we categorize existing datasets into unsupervised, weakly supervised, and supervised, as presented in Table 3. The examples of UVAD (e.g., Subway and UMN) and WsVAD datasets are illustrated in Fig. 7. Since such datasets have been extensively surveyed, we only present their details in Section 4 of Supplement. We provide a systematic WSVAD evaluation system in this section, categorizing existing metrics into accuracy-oriented and cost-oriented as well as introducing system-level performance indicators.
6.1. Conventional Evaluation
@@
The existing works typically evaluate the proposed method from two perspectives: detection accuracy and model cost. On the one hand, models are expected to detect noteworthy anomalies as accurately as possible. Considering the varying influences of false positives and false negatives, along with the highly imbalanced data, researchers have proposed multiple quantitative metrics such as Area Under the Receiver Operating Characteristic curve (AUROC), Area Under Precision-Recall curve (AUPR), and Detection Rate, to assess model accuracy. Additionally, anomaly score curves are commonly used to qualitatively demonstrate the sensitivity to abnormal intervals, while prediction error maps are widely employed in reconstruction-based UVAD methods to visualize the performance of spatial localization. On the other hand, metrics like parameter size and inference speed determine whether the model can be deployed on resource-constrained devices. Thus, we categorize conventional metrics based on the orientation and evaluation dimension, as illustrated in Fig. 8.
6.1.1. Accuracy-oriented metrics
@@
Due to the highly imbalanced nature of positive and negative samples in NSVAD, some common evaluation metrics for classification tasks, such as accuracy, are not applicable. For example, a model biased towards outputting label 0 thus missing anomalous events, would be incorrectly evaluated as good under such metrics. Using a single threshold to simply assess a model’s ability to differentiate between normal and abnormal patterns is not wise. For instance, with a threshold of 0.5, a model that consistently outputs scores slightly below 0.5 for regular events and slightly above 0.5 for anomalies would be considered perfect because it would show optimal performance across various metrics. However, such a model may not have learned the inherent differences between normal and abnormal patterns well, resulting in minimal score gaps between the two, which could lead to failure in detecting subtle anomalies and normal instances with data bias in complex scenarios. Therefore, researchers have introduced Receiver Operating Characteristic (ROC) curves, which measure VAD models more comprehensively by selecting multiple thresholds. Specifically, this curve plots the True Negative Rate (TNR) against the True Positive Rate (TPR) at various thresholds. The Area Under the ROC curve (AUROC) has been the most widely used VAD evaluation metric. An ideal AUROC value of 1 indicates a model that outputs a score of 0 for all negative samples and 1 for all positive samples, aligning with our expectations. Considering that TN usually exceeds TP, researchers argue that Average Precision, i.e., the Area Under the Precision-Recall (AUPR) curve, is more suitable for evaluating anomaly detection tasks. The PR curve depicts precision and recall (i.e., TPR) at specific thresholds. The point on this curve where Precision equals Recall is the balance point. Currently, multimodal anomaly detection models primarily use AP for quantitative evaluation. An anomaly score curve is commonly used to intuitively demonstrate the model’s response to anomalous events, presenting the temporal localization capability. In contrast, prediction error maps are often used to assess UVAD model’s spatial localization capability.
6.1.2. Cost-oriented metrics